# When isometry and equivalence for skew constacyclic codes coincide

### v1

@@ Abstract.
We show that the notions of $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence introduced by Ou-azzou et al coincide for most skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$. To prove this, we show that all Hamming-weight-preserving homomorphisms between their ambient algebras must have degree one when those algebras are nonassociative. We work in the general setting of commutative base rings $S$. As a consequence, we propose new definitions of equivalence and isometry of skew constacyclic codes that exactly capture all Hamming-preserving isomorphisms, and lead to tighter classifications. In the process we determine homomorphisms between nonassociative Petit algebras, prioritizing the algebras $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, which give rise to skew constacyclic codes.
Key words and phrases:
skew cyclic codes, skew polycyclic codes, quantum error-correcting codes, nonassociative algebra
@@ 1. Introduction
Petit algebras are nonassociative algebras that, when defined over finite fields or rings, are increasingly important for constructing and classifying skew polycyclic codes, as there is a one-one correspondence between skew polycyclic codes and principal left ideals in Petit algebras [Pum17]. Two such codes over a finite field are isometric if there is a Hamming-weight preserving isomorphism between them that comes from a monomial isomorphism between their corresponding Petit algebras viewed as algebras over the prime field, where the Hamming weight of a polynomial $p(t)$ is defined as the number of nonzero coefficients of $p(t)$. Such an isomorphism between Petit algebras, also called an isometry, thus gives a correspondence between their sets of associated codes that preserves code rate and Hamming distance as well as length and dimension [Pum25].
@@
There has been significant work to understand these isometries. In [OaHA25, OaNA25], the authors consider a subclass of such isometries (namely, those monomial isomorphisms that act as the identity on the finite field $\mathbb{F}_{q}$) to propose the notions of $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry between skew constacyclic codes (and of $(\mathbf{n},\sigma)$-equivalence for skew polycyclic codes) over $\mathbb{F}_{q}$, generalizing [CDL14]. The same subclass of isometries was used, but between two associative algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$, to define equivalences of skew constacyclic codes in [BBB21] and later again in [LMn25]. Skew constacyclic codes are the skew polycyclic codes obtained from Petit algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ (notation as in Section 2) for $S$ a unital commutative (associative) ring and $a\in S^{\times}$.
@@
In this paper we prove that these notions coincide for all classes of skew constacyclic codes over $S$ arising from proper nonassociative Petit algebras, a result that was only previously known for when the order $\mathbf{m}$ of $\sigma$ satisfied $\mathbf{m}\geq\mathbf{n}-1$ and when $S=K$ was a field [BP18].
@@
The motivation for the present work was two-fold: firstly, we wanted to understand what turned out to be an error in [OaNA25, Theorem 4], which suggested the existence of isometries of higher degree for all skew constacyclic codes. Searching for these led us to discover a rather startling result.
@@
To state it succinctly, note that we say that the algebra $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is proper nonassociative when it is not an associative algebra; by [Pet67, (15)] this occurs exactly when $\sigma(a)\neq a$ or $\mathbf{m}\nmid\mathbf{n}$.
@@
Let $K$ be a field with $\sigma\in\operatorname{Aut}(K)$ of finite order $\mathbf{m}$, and let $a_{1},a_{2}\in K^{\times}$. If the Petit algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a_{i})$ are proper nonassociative, then all nonzero Hamming weight preserving homomorphisms $G$ between them such that $G|_{K}$ commutes with $\sigma$ are monomial of degree one, that is, satisfy $G(t)=\alpha t$ for some $\alpha\in K^{\times}$.
In particular, the theorem implies that these algebras admit no nonzero Hamming weight preserving homomorphisms of the form $G(t)=\alpha t^{k}$ for any $\alpha\in K^{\times}$ and $1<k<\mathbf{n}$.
We attack this question with a novel approach, and in an important greater generality, by first characterizing the power-associative monomials $\alpha t^{k}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, where now we allow $S$ to be any commutative unital (associative) ring, and $\alpha,b\in S^{\times}$. The answer is quite elegant.
@@
The monomial $\alpha t^{k}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, with $\alpha\in S$ not a zero divisor and $1\leq k<\mathbf{n}$, is power-associative if and only if
(1.1) $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b).$
if we assume that $(\alpha t^{k})^{r}(\alpha t^{k})=(\alpha t^{k})(\alpha t^{k})^{r}$ for the least $r>0$ such that $rk\geq{\mathbf{n}}$.
The proof turns on a careful combinatorial argument which occupies most of Section 2. As a corollary, we give a classification of all monomial homomorphisms $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ for which the leading coefficient of $G(t)$ is not a zero divisor and for which the restriction of $G$ to $S$ is an automorphism commuting with $\sigma$ (Corollary 3.4).
@@
We prove in Corollary 3.4 that if $k>1$ and $\tau$ is an automorphism of $S$ commuting with the order-$\mathbf{m}$ automorphism $\sigma$, then being a homomorphism $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ given by $G(a)=\tau(a)$ and $G(t)=\alpha t^{k}$ entails $\mathbf{m}|\mathbf{n}$, $k\equiv 1\mod\mathbf{m}$ and thus $\sigma(b)=b$, whence $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative.
@@
We then leverage this to prove in Corollary 4.5 that if $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is proper nonassociative, then in fact there are no nonzero monomial homomorphisms $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ of degree other than one. Thus $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence as introduced in [OaNA25] coincide for classes of skew constacyclic codes that are associated with proper nonassociative Petit algebras.
@@
We then explain how our results lead to the failure of [OaNA25, Theorem 4] in Section 5, providing explicit counterexamples. We also recalculate the number of $(\mathbf{n},\sigma)$-equivalence classes of $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$, for $K$ a finite field, in Theorem 5.6.
@@
Moreover, in Section 5 we go one to define more general notions of equivalence and isometry that perfectly capture the set of isomorphisms that will preserve Hamming distance. This offers a tighter classification of equivalence and isometry classes of codes (see Definition 5.1). The advantage of this perspective is clearly demonstrated in Example 5.8). For associative algebras, a variation of [OaNA25, Theorem 4] can be proved (Theorem 5.9). In particular, the existence of a monomial isomorphism of degree $k>1$ in the associative case gives rise to monomial isomorphisms of degree one with another algebra.
Section 6 addresses our second motivation for this work which is to eventually extend our parametrization of proper nonassociative Petit division algebras of small dimension that covered the case that $\mathbf{n}=\mathbf{m}$ is prime and $f(t)=t^{\mathbf{n}}-a$ [NP25]. To extend these results to any $\mathbf{n}$ and $\mathbf{m}$ we need to understand all possible isomorphisms between two division algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$, not only the monomial ones.
@@
These division algebras are in one one correspondence with maximum rank distance codes. Any parametrization will thus yield parametrizations of the corresponding codes. For $\mathbf{n}=\mathbf{m}$, the division algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ have just been used in a variation of Learning with Errors, a lattice based approach to postquantum cryptography [ML25].
When $S$ is an integral domain, we prove that polynomial, i.e., non monomial, morphisms between properly nonassociative algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ do not occur, subject to a technical hypothesis we denote ($\star$6.3).
As a by product, we obtain from Theorem 6.1 that for integral domains we obtain that if $a,b\in S^{\times}$, and $\mathbf{m}\nmid\mathbf{n}$, then any nonzero homomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, must be monomial (Corollary 6.2).
@@
Skew constacyclic and skew polycyclic codes over finite rings and mixed alphabets are currently among the most studied kinds of linear codes, for their applications to constructions of quantum error-correcting codes (such as via Gray maps and the CSS construction [BDU+20, VPSI23]), with many papers appearing on this topic in recent years, including for example [BP25, PIP22, YSP23, PPI25, RSG24]. Our current work serves to improve the de-duplication of codes by classifying all isometries between them, which will in turn make the search for good quantum error-correcting codes more efficient.
In particular, when considering the CSS construction for quantum error-correcting codes, one needs to concentrate on codes $C$ such that the dual code $C^{\perp}$ satisfies $C^{\perp}\subset C$ as explained in [DBU+21]. Thus an important further extension of our work is to consider when our isometries preserve some suitable notion of duality for these codes.
2. Petit algebras and power associativity
@@
Let $S$ be a commutative, associative unital ring and suppose $\sigma\in\operatorname{Aut}(S)$ has finite order $\mathbf{m}$. Important examples in applications to coding theory are $S=K$ a finite field, or $S$ a chain ring (also called a finite local ring) or more generally a Galois ring (see, for example, [Pum17, §5.1]). In both these examples, every non-invertible element is a zero divisor.
@@
Write $S_{0}=\{s\in S:\sigma(s)=s\}$ for the subring fixed by $\sigma$. If $\tau\in\operatorname{Aut}(S)$ is an arbitrary automorphism then we write $S_{0}^{\tau}=\{s\in S_{0}:\tau(s)=s\}$ for the common fixed subring under $\tau$ and $\sigma$. We define $K_{0},K_{0}^{\tau}$ analogously.
@@
The ring $R=S[t;\sigma]$ is a noncommutative polynomial ring, with addition the usual addition of polynomials, and with multiplication defined by the relation
@@
for all $a\in S$ and all $j\in\mathbb{N}$. It is an algebra over $S_{0}$.
Let $f(t)=t^{\mathbf{n}}-f_{0}(t)\in S[t;\sigma]$ be a monic polynomial of degree $\mathbf{n}\geq 1$. For all $g(t)\in S[t;\sigma]$ of degree $\ell\geq\mathbf{n}$, there exist uniquely determined $r(t),q(t)\in S[t;\sigma]$ with ${\rm deg}(r)<{\rm deg}(f)$, such that $g(t)=q(t)f(t)+r(t)$ [Pum17, Proposition 1]. Let $r:=g\,{\rm mod}_{r}f$ denote the remainder of this right division of $g$ by $f$. Since the remainders are uniquely determined, we may identify
$S[t;\sigma]/S[t;\sigma]f(t)=\{\sum_{i=0}^{\mathbf{n}-1}b_{i}t^{i}:b_{i}\in S\}$
as left $S[t;\sigma]$-modules. In fact, this quotient also carries the structure of a (potentially) nonassociative unital ring that is an algebra over the fixed ring $S_{0}$, with multiplication of any two polynomials of degree less than $\mathbf{n}$ defined by
$g\circ h=gh\,\,{\rm mod}_{r}f.$

### v3

@@ Abstract.
We work in the setting of linear skew constacyclic codes over a commutative base ring $S$. We show that the notions of $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence introduced by Ou-azzou et al coincide for most skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$. To prove this, we show that all Hamming-weight preserving isomorphisms between their ambient rings which extend some automorphism $\tau$ of $S$ that commutes with $\sigma$ must have degree one, when those rings are not associative. In the process we determine isomorphisms between their nonassociative ambient rings, the Petit rings $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, which give rise to skew constacyclic codes. As a consequence, we propose new definitions of equivalence and isometry of skew constacyclic codes that exactly capture all Hamming-weight preserving isomorphisms between the ambient rings of skew constacyclic codes which extend $\tau\in\operatorname{Aut}(S)$ that commute with $\sigma$, and lead to tighter classifications.
@@ 1. Introduction
Let $S$ be a unital commutative (associative) ring, $a\in S$ invertible and $\sigma\in\operatorname{Aut}(S)$. Then a linear code $C\subset S^{\mathbf{n}}$ is called a skew $(\sigma,a)$-constacyclic code, if for every codeword $(c_{0},\dots,c_{\mathbf{n}-1})\in C$, the “shifted” codeword $(\sigma(c_{\mathbf{n}-1})a,\sigma(c_{0}),\dots,\sigma(c_{\mathbf{n}-2}))$ also lies in $C$. More generally, a linear code is a skew polycyclic code, or more precisely called a skew $(f,\sigma)$-polycyclic code, if every codeword satisfies a similar but more elaborate shift, this time determined by the coefficients of a pre-chosen reducible skew polynomial $f(t)\in S[t;\sigma]$.
Petit rings are nonassociative unital rings that can be considered as the ambient rings of these skew polycyclic codes, as there is a one-one correspondence between a skew polycyclic code and a principal left ideal in a Petit ring [Pum17, Pum25]. For instance, skew $(\sigma,a)$-constacyclic codes correspond to the left principal ideals in their nonassociative ambient Petit ring $\mathbb{S}_{(t^{\mathbf{n}}-a)}$, whose underlying module is $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ (notation as in Section 2.1). Left multiplication of any element in a left principal ideal of the ambient (albeit potentially nonassociative) Petit ring by the indeterminate $t$ then creates the shift employed in the definition of a skew polycyclic code. Thus while skew $(\sigma,a)$-constacyclic codes are often viewed only as $S[t;\sigma]$-submodules of $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, using the perspective of nonassociative Petit rings significantly enriches the algebraic tools available for classification, as this paper illustrates.
@@
The Hamming-weight of a codeword $c=(c_{0},\dots,c_{\mathbf{n}-1})$ is defined as the number of nonzero coefficients of $c$, the Hamming-weight of a polynomial $p(t)$ is defined as the number of nonzero coefficients of $p(t)$. Two skew polycyclic codes are called isometric if there is a Hamming-weight preserving isomorphism between them that comes from a monomial isomorphism between their corresponding Petit rings, that means the isomorphism, denoted here $G_{\tau,\alpha,k}$, maps $t$ to $\alpha t^{k}$ for some positive integer $k$ and invertible $\alpha\in S$. Such an isomorphism between Petit rings, also called an isometry, preserves the Hamming-weight of a polynomial and, thus gives a correspondence between the sets of associated codes that preserves code rate and Hamming distance as well as length and dimension [Pum25].
@@
There has been significant work to understand the isometries of skew polycyclic codes. In [OaHA25, OaNA25], the authors consider a subclass of such isometries (namely, those monomial isomorphisms that act as the identity on the finite field $\mathbb{F}_{q}$) to propose the notions of $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry between skew constacyclic codes (and of $(\mathbf{n},\sigma)$-equivalence for skew polycyclic codes) over $\mathbb{F}_{q}$, generalizing [CDL14]. The same subclass of isometries was used, but between the algebras $\mathbb{F}_{q}[t;\sigma]/\mathbb{F}_{q}[t;\sigma](t^{\mathbf{n}}-a)$ and $\mathbb{F}_{q}[t;\sigma]/\mathbb{F}_{q}[t;\sigma](t^{\mathbf{n}}-b)$, to define equivalences of skew constacyclic codes in [BBB21] and later again in [LMn25]. However, this latter paper only considered a restricted case: where $\mathbf{n}$ is a multiple of the order of $\sigma$ and $a$ fixed by $\sigma$; in this case the ambient ring is associative.
@@
In this paper we generalize and prove that the notions of $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry coincide for all classes of skew constacyclic codes over $S$ whose ambient Petit rings are not associative. This result was only previously known for when the order $\mathbf{m}$ of $\sigma$ satisfied $\mathbf{m}\geq\mathbf{n}-1$ and when $S=K$ was a field [BP18].
Readers interested in codes can simply skip the technical Sections 3, 4 and 6 that prove the requisite properties of isometries of Petit rings, and proceed to Section 5 directly, which applies these results to skew constacyclic codes.
@@
The motivation for the present work was two-fold: first, we wanted to understand what turned out to be an error in [OaNA25, Theorem 4], which suggested the existence of isometries of higher degree for all skew constacyclic codes. Searching for these led us to discover a rather startling result.
@@
To state it succinctly, let $S=K$ be a finite field. Write $\mathbb{K}_{t^{\mathbf{n}}-a}$ for the Petit ring corresponding to $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$. It is said to be proper nonassociative when it is not an associative ring; by [Pet67, (15)] this occurs exactly when $\sigma(a)\neq a$ or $\mathbf{m}\nmid\mathbf{n}$.
@@
Let $K$ be a field with $\sigma\in\operatorname{Aut}(K)$ of finite order $\mathbf{m}$, and let $a_{1},a_{2}\in K^{\times}$. If the Petit rings $\mathbb{K}_{t^{\mathbf{n}}-a_{i}}$ are proper nonassociative, then all nonzero Hamming-weight preserving homomorphisms $G$ between them such that $G|_{K}$ commutes with $\sigma$ are monomial of degree one, that is, satisfy $G(t)=\alpha t$ for some $\alpha\in K^{\times}$.
In particular, the theorem implies that these rings admit no nonzero Hamming-weight preserving homomorphisms of the form $G(t)=\alpha t^{k}$ for any $\alpha\in K^{\times}$ and $1<k<\mathbf{n}$.
Our approach is novel, and applies in an important greater generality, where the field $K$ is replaced by any commutative unital (associative) ring $S$ (such as a Galois ring, for example). The first step is to characterize the power-associative monomials $\alpha t^{k}\in\mathbb{S}_{(t^{\mathbf{n}}-b)}$, where $\alpha,b\in S^{\times}$. The answer is quite elegant and presented in Section 3.
@@
Suppose $\alpha\in S$ is not a zero divisor and $1\leq k<\mathbf{n}$. The monomial $\alpha t^{k}\in\mathbb{S}_{(t^{\mathbf{n}}-b)}$ is power-associative if and only if
(1.1) $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)\in\mathbb{S}_{t^{\mathbf{n}}-b}.$
The proof turns on a careful combinatorial argument which occupies most of Section 3.
@@
As a corollary, we give a classification of all monomial homomorphisms $G:S[t;\sigma]\to\mathbb{S}_{t^{\mathbf{n}}-b}$ of the form $G|_{S}=\tau$ and $G(t)=\alpha t^{k}$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, $\alpha\in S^{\times}$ and $1\leq k<\mathbf{n}$ (Corollary 3.4). In particular, if such a homomorphism with $k>1$ exists, then this entails $\mathbf{m}|\mathbf{n}$, $k\equiv 1\mod\mathbf{m}$ and thus $\sigma(b)=b$, whence $\mathbb{S}_{t^{\mathbf{n}}-b}$ is associative.
@@
We then leverage this to prove in Corollary 4.5 that if $\mathbb{S}_{t^{\mathbf{n}}-b}$ is proper nonassociative, then in fact there are no nonzero monomial homomorphisms $G:\mathbb{S}_{f}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ of degree other than one. Thus $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence as introduced in [OaNA25] coincide for classes of skew constacyclic codes that are arise from proper nonassociative Petit ambient rings.
@@
We then explain how our results lead to the failure of [OaNA25, Theorem 4] in Section 5, providing explicit counterexamples. We also recalculate the number of $(\mathbf{n},\sigma)$-equivalence classes of skew constacyclic codes with ambient algebras $\mathbb{K}_{t^{\mathbf{n}}-a}$, for $K$ a finite field, in Theorem 5.9.
@@
In Section 5 we go one to define more general notions of equivalence and isometry that perfectly capture the set of isomorphisms that will preserve Hamming distance. This offers a tighter classification of equivalence and isometry classes of codes (see Definition 5.1). The advantage of this perspective is clearly demonstrated in Example 5.11. For associative algebras, we can prove a variation of [OaNA25, Theorem 4] (Theorem 5.12). In particular, the existence of a monomial isomorphism of degree $k>1$ between $\mathbb{S}_{t^{\mathbf{n}}-a}$ and $\mathbb{S}_{t^{\mathbf{n}}-b}$ in the associative case gives rise to monomial isomorphisms of degree one from $\mathbb{S}_{t^{\mathbf{n}}-a}$ to $\mathbb{S}_{t^{\mathbf{n}}-b^{k}}$.
@@
Section 6 addresses our second motivation for this work which is to eventually extend our parametrization of proper nonassociative Petit division algebras begun in [NP25]. In that paper, we solved the case that $\mathbf{n}=\mathbf{m}$ is prime and $f(t)=t^{\mathbf{n}}-a$. These division algebras, in turn, are in one one correspondence with maximum rank distance codes. Any parametrization will thus yield parametrizations of the corresponding codes. For $\mathbf{n}=\mathbf{m}$, the division algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ have just been used in a variation of Learning with Errors, a lattice based approach to postquantum cryptography [MenLin25].
Crucially, to extend the result in [NP25] to any $\mathbf{n}$ and $\mathbf{m}$ one needs to understand all possible isomorphisms between two division algebras $\mathbb{K}_{t^{\mathbf{n}}-a}$ and $\mathbb{K}_{t^{\mathbf{n}}-b}$, not only the monomial ones. When $S$ is an integral domain, we prove that polynomial, i.e., non monomial, morphisms between proper nonassociative algebras of the form $\mathbb{S}_{t^{\mathbf{n}}-a}$ do not occur, subject to a technical hypothesis we denote ($\star$6.3). Consequently, it follows from Theorem 6.1 that for integral domains $S$, $a,b\in S^{\times}$, and $\mathbf{m}\nmid\mathbf{n}$, every nonzero homomorphism $G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, must be monomial of degree one (Corollary 6.2).
@@
Skew constacyclic and skew polycyclic codes over finite rings and mixed alphabets are currently among the most studied kinds of linear codes. Our current work serves to improve the de-duplication of codes by classifying all isometries between them.
We point out that our results for isometries between the ambient Petit algebras also have consequences for skew polycyclic codes over finite fields with the rank metric, where we need $\alpha\in K^{\prime}$ for a suitable subfield of $K$ we consider the rank metric over. In particular, they will help understand the different MRD codes obtained from these [OaHA25, Proposition 3.13].
2. Preliminaries
@@
Let $S$ be a commutative, associative unital ring, $S^{\times}$ its set of invertible elements, and suppose $\sigma\in\operatorname{Aut}(S)$ has finite order $\mathbf{m}$. Important examples in applications to coding theory are $S=K$ a finite field, $S$ a Galois ring or more generally $S$ a chain ring (also called a finite local ring); see for example [Pum17, §5.1]). In all these examples, every non-invertible element is a zero divisor.
@@
Write $S_{0}=\{s\in S:\sigma(s)=s\}$ for the subring fixed by $\sigma$. If $\tau\in\operatorname{Aut}(S)$ is an automorphism commuting with $\sigma$ then we write $S_{0}^{\tau}=\{s\in S_{0}:\tau(s)=s\}$ for the common fixed subring under $\tau$ and $\sigma$.
@@
The ring $R=S[t;\sigma]$ is a unital associative noncommutative skew polynomial ring, with addition the usual addition of polynomials, and with multiplication defined by the relation
@@
for all $a\in S$ and all $j\in\mathbb{N}$.
Let $f(t)=t^{\mathbf{n}}-f_{0}(t)\in S[t;\sigma]$ be a monic polynomial of degree $\mathbf{n}\geq 1$. For all $g(t)\in S[t;\sigma]$ of degree $\ell\geq\mathbf{n}$, there exist uniquely determined $r(t),q(t)\in S[t;\sigma]$ with ${\rm deg}(r)<{\rm deg}(f)$, such that $g(t)=q(t)f(t)+r(t)$, e.g. see [Pum17, Proposition 1]. Since the remainders are uniquely determined, we may identify the additive left cosets of the left ideal $S[t;\sigma]f(t)$ with the set
$S[t;\sigma]/S[t;\sigma]f(t)=\{\sum_{i=0}^{\mathbf{n}-1}b_{i}t^{i}:b_{i}\in S\}=:S[t;\sigma]_{\mathbf{n}}.$
This is a familiar module over $S[t;\sigma]$, but a priori it is not clear that there is always a well-defined ring multiplication on this set of representatives $S[t;\sigma]_{\mathbf{n}}$ as well.
We call $f(t)\in R$ two-sided, if $S[t;\sigma]f(t)$ is a two-sided ideal. It is well-known that for two-sided skew polynomials $f(t)$ (e.g., $f(t)$ that lie in the center of $S[t;\sigma]$), we can factor out the two-sided ideal $S[t;\sigma]f(t)$ in the associative ring $S[t;\sigma]$ to obtain the associative quotient ring $S[t;\sigma]/S[t;\sigma]f(t)$ whose multiplication of any two polynomials of degree less than $\mathbf{n}$ is defined by
$g\circ h=gh\,\,{\rm mod}_{r}f,$
where $g\,{\rm mod}_{r}f$ denotes the remainder of right division of $g$ by $f$.
2.1. Petit rings are not quotient rings
Even when $f(t)\in S[t;\sigma]$ is not two-sided, we define an operation on $S[t;\sigma]_{\mathbf{n}}$ by the rule
$g\circ h=gh\,\,{\rm mod}_{r}f$
and this endows $S[t;\sigma]_{\mathbf{n}}$ with the structure of a unital nonassociative ring. We call the unital ring $(S[t;\sigma]_{\mathbf{n}},\circ)$ defined in this way a Petit ring and note that it can also be viewed as an algebra over $S$ (or over any subring of $S$). Throughout this paper we write $\mathbb{S}_{f}=(S[t;\sigma]_{\mathbf{n}},\circ)$ for this nonassociative ring following current literature, but often also simply $S[t;\sigma]/S[t;\sigma]f(t)$ if we want to make clear which $S$ and which $\sigma$ are involved. It is then clear from the context that we talk about nonassociative Petit rings. When $S=K$ is a field we may write $\mathbb{K}_{f}$ for $K[t;\sigma]/K[t;\sigma]f(t)$.
This multiplication is not associative in general. We know that $\mathbb{S}_{f}=S[t;\sigma]/S[t;\sigma]f(t)$ is an associative ring exactly when $S[t;\sigma]f(t)$ is a two-sided ideal [Pet67]. For example, this happens when $S=K$, $f(t)=t^{\mathbf{n}}-\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\in S_{0}[t]$ and $\sigma^{\mathbf{n}-i}=\operatorname{id}$ for all those $i$ where $a_{i}\not=0$ ([Pet67, (15)]. In that setting, the two-sided elements $f(t)$ are all of the form $ac(t)t^{\ell}$, where $a\in K^{\times}$, where $c(t)$ lies in the center of $K[t;\sigma]$ and $\ell\geq 0$ is an integer [Jac1996, Theorem 1.1.22].
The definition of this nonassociative ring structure on the representatives of $S[t;\sigma]/S[t;\sigma]f(t)$ can be seen as the canonical generalization of the classical quotient ring.